OpenStudy (anonymous):

which of the following is the solution to the equation 8(y + 3) = 2y ?

OpenStudy (anonymous):

which of the following is the solution to the equation 8^(y + 3) = 2^y ?

OpenStudy (bahrom7893):

2^(3(y+3))=2^y 3(y+3)=y

OpenStudy (bahrom7893):

3y+9=y 2y=-9 y=-9/2